OpenStudy (anonymous):

What is the y-value of cos(x) when x = 90°?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello

OpenStudy (misty1212):

are you looking for \[\cos(90^\circ)\]?

OpenStudy (anonymous):

is that the answer?

OpenStudy (misty1212):

lol no dear, i think that is what you are trying to find, the cosine of ninety degrees

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont understand cosine at all

OpenStudy (anonymous):

i dont

OpenStudy (anonymous):

so 1 is the answer?

OpenStudy (misty1212):

now you do

OpenStudy (misty1212):

look at the unit circle on the last page of the cheat sheet find the angle \(90^\circ\) and look at the coordinates of that point you will see \((0,1)\) the first coordinate is cosine, the second is sine

OpenStudy (misty1212):

so \[\cos(90)=0,\sin(90)=1\]

OpenStudy (anonymous):

ah i see

OpenStudy (misty1212):

i made a mistake earlier and said \(\cos(90)=1\) what was wrong it is \(\cos(90)=0\)

OpenStudy (anonymous):

its looking for the y value, so the answer is still 1?